# FedQUIT: On-Device Federated Unlearning via a Quasi-Competent Virtual Teacher

### v2

@@ 15 Further Experimental Results
Algorithm Rounds ($\downarrow$) CR ($\uparrow$) Test Acc. Forget Acc. MIA [29] Original 75.03 ±0.00 84.25 ±5.63 77.03 ±8.57 Retrained 73.30 ±0.78 57.80 ±5.29 48.59 ±4.96 PGA [12] 6.9 ±4.28 7.2$\times$ 73.60 ±0.36 62.15 (4.35 ±3.28) 53.57 (4.98 ±3.67) FedQUIT 8.78 ±2.54 5.7$\times$ 73.42 ±0.89 58.50 (1.86 ±1.46) 48.86 (2.56 ±2.13)
@@
Setting KD Layer $v$ Rounds ($\downarrow$) CR ($\uparrow$) Test Acc. Forget Acc. MIA [29] MIA [36] CIFAR-100, IID, $E$=1 Logit 0 10.0±1.4 20.0$\times$ 58.70 ±0.4 59.43 (1.4±0.8) 60.71 (5.1±1.6) 55.70 (6.8±0.8) Logit min 10.1±1.73 58.59±0.58 59.66 (1.70±1.01) 61.96 (6.32±1.78) 55.92 (7.08±0.86) Softmax $1/C$ 8.1±1.66 58.61±0.40 67.23 (9.26±0.85) 66.63 (11.01±1.52) 61.72 (12.88±1.10) Softmax 0 8.3±1.34 58.57±0.43 67.95 (9.98±0.91) 66.34 (10.73±1.00) 61.98 (13.14±1.08) CIFAR-100, Non-IID, $E$=1 Logit 0 3.7±2.9 54.1$\times$ 52.00±0.5 34.60 (3.5±2.4) 50.74 (5.5±2.8) 33.93 (3.8±2.8) Logit min 3.2±1.87 51.66±1.23 34.10 (3.05±2.05) 50.10 (6.08±3.32) 33.54 (3.11±2.29) Softmax $1/C$ 3.4±3.75 51.55±1.62 40.31 (6.85±3.83) 53.86 (9.8±4.11) 38.60 (6.59±3.58) Softmax 0 2.4±1.9 51.57±1.32 39.09 (5.97±3.52) 52.31 (8.39±4.62) 36.81 (5.92±2.91) CIFAR-100, Non-IID, $E$=10 Logit 0 1.7±0.5 117.6$\times$ 48.30 ±0.4 27.93 (3.4±2.4) 45.90 (4.7±2.8) 29.90 (4.1±2.9) Logit min 2.00±1.00 48.81±0.89 29.45 (4.14±3.24) 47.33 (5.29±4.70) 31.04 (3.90±3.22) Softmax $1/C$ 0.71±1.67 48.85±0.95 38.82 (7.43±4.85) 52.14 (10.02±5.18) 37.98 (6.88±5.22) Softmax 0 2.11±0.60 49.00±0.87 40.20 (8.81±5.23) 54.01 (11.78±5.57) 39.81 (8.72±5.73)
Table 10: FedQUIT performance across different KD layers (logits or softmax output) and values of $v$ used in 3 (FedQUIT-Logit) and 4 (FedQUIT-Softmax). The table reports the number of communication rounds and forgetting accuracy across IID and Non-IID settings, including scenarios with $E$=10.
@@
Algorithm Round Test Acc. Forget Acc. MIA [29] MIA [36] Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0 ±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0 ±5.2 Natural 1 53.56 (2.56±1.01) 54.28 (20.78±7.23) 67.32 (23.32±7.86) 52.48 (20.48±7.33) Natural 2 53.41 (2.41±1.63) 52.05 (18.55±7.08) 64.93 (20.93±7.70) 49.91 (17.91±7.33) Natural 4 53.29 (2.29±1.45) 50.49 (16.99±6.06) 63.11 (19.11±7.19) 48.44 (16.44±6.26) Natural 16 53.31 (2.31±1.30) 45.74 (12.24±4.84) 57.86 (13.86±5.73) 43.16 (11.16±4.52) Natural 20 53.67 (2.67±1.35) 45.37 (11.87±4.84) 56.97 (12.97±5.99) 42.18 (10.18±4.80)

### v3

@@ 15 Further Experimental Results
[…]
@@
[…]
Table 10: FedQUIT performance across different KD layers (logits or softmax output) and values of $v$ used in 3 (FedQUIT-Logit) and 4 (FedQUIT-Softmax). The table reports the number of communication rounds and forgetting accuracy across IID and Non-IID settings, including scenarios with EE=10.
@@
Algorithm Round Test Acc. Forget Acc. MIA [song2021systematic] MIA [yeom2018privacy] Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0 ±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0 ±5.2 Natural 1 53.56 (2.56±1.01) 54.28 (20.78±7.23) 67.32 (23.32±7.86) 52.48 (20.48±7.33) Natural 2 53.41 (2.41±1.63) 52.05 (18.55±7.08) 64.93 (20.93±7.70) 49.91 (17.91±7.33) Natural 4 53.29 (2.29±1.45) 50.49 (16.99±6.06) 63.11 (19.11±7.19) 48.44 (16.44±6.26) Natural 16 53.31 (2.31±1.30) 45.74 (12.24±4.84) 57.86 (13.86±5.73) 43.16 (11.16±4.52) Natural 20 53.67 (2.67±1.35) 45.37 (11.87±4.84) 56.97 (12.97±5.99) 42.18 (10.18±4.80)